OpenStudy (anonymous):

In what direction is the derivative of f( x,y)= xy+y^2 at P(3, 2) equal to zero?

OpenStudy (abb0t):

Where you given a direction vector?

OpenStudy (amistre64):

doesnt the normal at a given point on a surface point towards the highest point?

OpenStudy (amistre64):

not the normal, but the gradient .... fx = 1, fy = 1+2y <fx,fy> = <1,5>

OpenStudy (amistre64):

the dot product of gradf and the direction vector comes to mind, and they want the to be zero, so the direction vector must be the perp .... <1,5> perps to <5,-1> or <-5,1>

zepdrix (zepdrix):

I'm trying to remember how to do this... "In what direction is the derivative of.... ....equal to zero" So we want to set up a directional derivative, and set it equal to zero, and solve for the vector. \[\huge Dir_u f(x,y)\qquad = \qquad \nabla f \cdot \;\hat u\] \[\huge Dir_u f(x,y)\qquad = \qquad <f_x,f_y>\cdot \;\hat u=0\]This reads "The directional derivative of f, in the direction of u, equals ..." Hmm let's see. Taking the partials gives us,\[\large Dir_u f(x,y) \qquad = \qquad <y+y^2,\quad x+2y>\cdot \;\hat u\] Evaluated at the point (3,2), \[\large Dir_u f(3,2) \qquad = \qquad <2+2^2,\quad 3+2\cdot2>\cdot \;\hat u\]And then setting this equal to zero,\[\large <6,\quad 7>\cdot \;\hat u \qquad = 0\] And this is only true when u is perpendicular to our vector <6,7>. I think @amistre64 had the right idea. I just think there was a little miscalculation in the partials.

zepdrix (zepdrix):

Woops, I messed up my partials also... The second part of \(f_x\) was a constant, which should have become 0. So \(f_x\) should give us \(y\), not \(y+y^2\).

OpenStudy (amistre64):

yeah, i messed up on the fx fy parts :) the rest of my recollection seems to have been sound tho ;) thnks